# Context-Parametric Inversion:Why Instruction Finetuning May Not Actually Improve Context Reliance

### v1

@@ Abstract
Large language models are instruction-finetuned to enhance their ability to follow user instructions and process the input context. However, even state-of-the-art models often struggle to follow the instruction, especially when the input context is not aligned with the model’s parametric knowledge. This manifests as various failures, such as hallucinations where the responses are outdated, biased or contain unverified facts. In this work, we try to understand the underlying reason for this poor context reliance, especially after instruction tuning. We observe an intriguing phenomenon: during instruction tuning, the context reliance initially increases as expected, but then gradually decreases as instruction finetuning progresses. We call this phenomenon context-parametric inversion and observe it across multiple general purpose instruction tuning datasets like TULU, Alpaca and Ultrachat, as well as model families such as Llama, Mistral and Pythia. In a simple theoretical setup, we isolate why context-parametric inversion occurs along the gradient descent trajectory of instruction finetuning. We tie this phenomena to examples in the instruction finetuning data mixture where the input context provides information that is already present in the model’s parametric knowledge. Our analysis suggests natural mitigation strategies that provide gains across models and datasets. We hope that our work serves as a starting point in addressing this failure mode in a staple part of LLM training.
@@ 1 Introduction
Large language models (LLMs) have unlocked remarkable capabilities across a wide range of tasks, excelling in understanding, generation, and reasoning. These advancements are driven by heavy pretraining on vast amounts of rich text data, followed by instruction finetuning to teach models to better comprehend and follow user instructions. However, one key challenge is ensuring that models prioritize the input context over their vast pretrained knowledge store. Overreliance on parametric knowledge can lead to responses that are outdated, biased, missing minority perspectives, or can cause hallucinations by introducing unverified facts into model’s responses  (Qiu et al., 2023; Adlakha et al., 2024; Longpre et al., 2022; Shi et al., 2023).
Chatbots particularly undergo “instruction finetuning” to better assist users with tasks and queries. Oftentimes, users supplement queries with an input context — an image, file, or document that provides the information needed to complete the task. For instance, users may request a summary of a recent news article or the next line of an incomplete block of code. However, while the instruction remains static, the context can rapidly diverge from the facts and patterns the models learned during pretraining as time passes or as they are deployed across different regions. Under such circumstances, state-of-the-art models have a tendency to answer a user query using their parametric knowledge over the input context. This been commonly studied under the moniker of knowledge conflicts (Shi et al., 2023; Jin et al., 2024a) and this problem intensifies with scale — larger models store more facts and thus face more conflicts (Longpre et al., 2022). Ideally, models should effectively parse the user input context when available and, otherwise, provide their best inference.
Figure 1: Context-Parametric Inversion (a) While instruction tuning is expected to improve context reliance, we observe that dependency on counterfactual conflicts surprisingly drops, after an initial increase. (b) Instruction finetuning data contain examples that may be answered using the context or the model’s parametric knowledge. These examples are harmful for context reliance.
@@
We strive to understand the underlying reason for this poor context reliance, especially even after instruction finetuning. In this work, we uncover an intriguing finding: during the process of instruction tuning, the context reliance on examples where the context is in conflict with the parametric knowledge, decreases with instruction finetuning, despite an initial expected increase. On the other hand, the performance on standard benchmarks (e.g., MMLU, GSM8k, SQuAD) keeps on increasing far after this drop. For example, as shown in Figure 1(a), the context-reliance of Llama2-7B (as measured on context-parametric conflict datasets (§ 3.2)) increases from 30% to 60% initially with instruction tuning. However, it start dropping as the finetuning progresses further. We observe this behavior on multiple instruction tuning datasets like TULU, Alpaca or UltraChat and across multiple model families like Llama, Pythia and Mistral. We call this phenomenon context-parametric inversion.
In principle, instruction tuning should improve a model’s ability to follow user instructions, including whatever information is provided in the context, and in fact we do see an initial increase. The subsequent decrease is detrimental and unexpected — there is no data in the most common instruction tuning sets that explicitly discourage the use of contextual information. We perform controlled studies to first eliminate our initial hypotheses. We find that this decrease is not simply due to memorizing facts during the finetuning stage and such facts causing further conflict at test-time: model learns to rely on parametric knowledge broadly well outside the exact facts seen during finetuning. Instruction tuning datasets contain a mixture of points (Figure 1(b)), some of which pair user instructions with a context while others don’t. Interestingly, even when we finetune only on a context-based subset, this phenomenon continues to occur.
@@
Our experiments show that there are context-critical datapoints where context provides key information needed to answer a user query that the model does not know beforehand (Fig. 1(b)). Conversely, there are non-critical datapoints where context overlaps with the model’s pretrained knowledge, making both context and parametric knowledge predictive. In the early stages of training, context-critical points tend to have higher loss and therefore dominate the gradient signal, driving the model to focus on the context. However, as training progresses, the loss on context-critical points decreases, and the non-context-critical points dominate the gradient signal. At later stages, the model leverages its pretrained knowledge to further reduce the loss on these non-context-critical points without increasing the average loss. As we will explain, this switch between using the context and the parametric knowledge can generalize beyond finetuning examples. We formalize this dynamic in a one-layer transformer under natural simplifying assumptions when finetuning on a data mixture consisting of context-critical and non-context-critical points.
Finally, our analysis naturally lead us to some mitigation strategies based on data curation, data augmentation, and regularization techniques. We test these out on deep networks on real-world datasets and see that they offer some gains, showing that our theoretical insights do translate to real-world settings. However, as we discuss in § 6, these mitigation strategies offer limited gains and come with tradeoffs, but we hope they inspire future works in this direction. Overall, in this work, we demonstrate a broad failure in instruction tuning, where we show that the model actually tends to increasingly rely more on the parametric knowledge than the context, despite an initial increase in context reliance. To the best of our knowledge, we are the first to point out this deficiency with instruction tuned models. We hope the conceptual explanation and mitigation strategies serve as useful starting points in addressing this fundamental tension between context and parametric knowledge during instruction tuning of LLMs.
@@ 2 Related Works
LLMs preferring their parametric knowledge over the user input context has been extensively studied under “knowledge conflicts” (Longpre et al., 2022; Shi et al., 2023; Yuan et al., 2024; Zhou et al., 2023; Xie et al., 2024). These works mainly focus on improving the faithfulness to the context using training-free approaches or by finetuning on counterfactual datasets. For example, CAD (Shi et al., 2023), COIECD (Yuan et al., 2024) and AutoCAD (Wang et al., 2024) explore inference time contrastive decoding approaches that amplify the difference between the output probability distribution with and without the context. These methods provide limited gains, especially in instruction finetuned models (Wang et al., 2024)Zhou et al. (2023); Zhang & Choi (2024) explore various prompting strategies to bias the model’s behavior towards context.  Jin et al. (2024b) tries to build a mechanistic interpretation. On the other hand,  Longpre et al. (2022); Fang et al. (2024); Neeman et al. (2022); Li et al. (2022) explore finetuning with counterfactual augmented data to mitigate knowledge conflicts (In § 6 we show that gains are limited to domains similar to the augmented data). Many evaluation studies are conducted on contexts that are easy for the model to process (Chen et al., 2022; Tan et al., 2024), which may downplay the pervasiveness of knowledge conflicts. In contrast, we take a step back and identify and highlight the root cause specifically during instruction finetuning that exacerbates model’s parametric reliance. Our observations show that models could have had much better context reliance as shown by the initial increase but subsequent decline during instruction finetuning.
@@
Retrieval augmented generation (RAG), where chat assistant systems retrieve relevant context to supplement the model generation, is very related to our setting. In fact, instruction finetuning with examples with input-contexts is very similar in practice to RAG. RAG has many scenarios of conflicts both between the various external sources or between external sources and parametric knowledge (Jin et al., 2024a; Kortukov et al., 2024). These works similarly highlight a confirmation bias where models tends to follow the evidence that aligns with their pretraining knowledge.  Guu et al. (2020) incorporate a retriever during pretraining to improve context reliance, whereas  Lewis et al. (2021) incorporate a retriever during finetuning. Several works suggest that context-reliance may sometimes not be desirable, especially when the input context is noisy and irrelevant.  Zhang et al. (2024) propose a training procedure to ignore irrelevant contexts.
Instruction Finetuning:
Instruction finetuning (IFT) is done to improve models ability to comprehend user input and instructions (Ding et al., 2023b). Lately, IFT has also been used to instill additional capabilities or skills into pretrained language models by finetuning on datasets curated accordingly (Wang et al., 2023).  Biderman et al. (2024); Wang et al. (2022); Kotha et al. (2024); Luo et al. (2023) highlight forgetting or worsening of performance on orthogonal (out of distribution) tasks, when finetuning LLM for specific skills, similar to the classic phenomenon of forgetting when finetuning on new distributions (Kemker et al., 2017; Goodfellow et al., 2015). In contrast, in this work we show an unexpected drop in context-reliance with instruction finetuning, after an expected initial increase. This is intriguing, as instruction finetuning is a ubiquitous approach used to improve LLMs ability to comprehend user instruction and context-reliance.
@@
We begin by observing context-parametric inversion across different models and datasets, by tracking context-reliance across the instruction finetuning trajectory.
Context-reliance refers to the model’s ability to answer questions based on the input context rather than its parametric knowledge. We are interested in the hardest scenario where these two sources provide opposing information. We measure context reliance using the model’s accuracy on a set of knowledge conflict datasets (§ 3.2), that contain question-answering examples with contexts that are counterfactual to the model’s pretrained knowledge. We measure accuracy by entailment. Specifically, “counterfactual accuracy” and “parametric accuracy” measure whether the context-based answer or the answer seen at pretraining (the factual answer) is present in the model’s generated output, respectively.
@@ 3.1 Experiment Setup
We experiment using three open source large language models—Llama2-7B, Pythia6.9B, and Mistral7B. We finetune for up to 2 epochs on three common instruction tuning datasets— TULU (Wang et al., 2023), UltraChat (Ding et al., 2023a), and Alpaca (Taori et al., 2023). We track the progress of instruction finetuning based on the performance on four standard benchmarks: GSM8k (Cobbe et al., 2021) (math), MMLU (Hendrycks et al., 2021) (general fact recall), SQuAD (Rajpurkar et al., 2016) (reading comprehension), and ARC-Challenge (Clark et al., 2018) (reasoning). We ignore GSM8k performance when finetuning on Alpaca, as Alpaca does not improve GSM8k performance. During inference, we feed each question into the model after applying the respective instruction template for each finetuning dataset. We refer the reader to Appendix A.2 for additional details.
@@ 3.2 Knowledge Conflict Datasets
For our study, we first carefully design three knowledge conflict datasets to get an accurate measure of model’s context reliance. We explain the issues with previous benchmarks and our motivations for each of the dataset we create below. All datasets are available at https://github.com/locuslab/context-parametric-inversion. We refer the reader to Appendix A.5 for some examples.
@@
    Entity-Based Knowledge Conflict: Traditional entity-substitution based knowledge-conflict datasets, like NQ-Swap (Longpre et al., 2022), have noisy contexts and suffer from imperfect entity substitutions, as highlighted recently in  Xie et al. (2024). This happens because the entity substitution models (Honnibal & Montani, 2017) are not able to recognize and replace all the occurrences of factual answers in the input. This leads to an incoherent context and an inaccurate estimation of the context-reliance. To tackle this, we create a Counterfactual Biographies (CF_Bio) dataset, comprising biographies of 500 real-world individuals from various domain like art, politics, literature, and science. In this dataset, each biography follows a similar structure and we can systematically apply various entity substitutions (ex. substituting names, contribution, etc.) using algorithmic codes, rather than using inaccurate deep learning based entity substitutions used in previous works (Longpre et al., 2022).
@@
    Coherent Counterfactual Contexts: Recently  Xie et al. (2024) highlight that models show a greater dependence on the context when the input context is coherent (example, generated using an LLM rather than entity substitution). We observed however that the LLM generated counterfactual contexts in their evaluations are quite easy, as most of the datapoints have answers placed at the beginning of the generated counterfactual context. Hence, we create a synthetic Counterfactual World Facts (CF_World_Facts) dataset, containing 400 questions with alternative explanations about counterfactual world facts generated using ChatGPT. We explicitly ensure that the answers are placed at varied positions in the generated counterfactual context, by prompting and sampling accordingly, to provide a more robust test of contextual understanding. We refer the reader to Appendix A.5 for further details and examples.
@@
    Beyond Context-Based QA: The tension between context and parameteric reliance goes beyond question-answering. It also applies to instruction following that might result in answers that contradict with parametric knowledge or well-known behaviors. For example, “Write a phrase that ends in heavy. Absence makes the heart grow ”. While the instruction requires the answer to be the word “heavy”, the parametric knowledge would suggest “fonder”. To measure context reliance in these cases, we use the Memo Trap task from the inverse scaling benchmark (McKenzie et al., 2024), and refer to it as CF_Quotes.

### v3

# Context-Parametric Inversion:Why Instruction Finetuning Can Worsen Context Reliance

@@ Abstract
A standard practice when using large language models is for users to supplement their instruction with an input context containing new information for the model to process. However, models struggle to reliably follow the input context, especially when it conflicts with their parametric knowledge from pretraining. In-principle, one would expect models to adapt to the user context better after instruction finetuning, particularly when handling knowledge conflicts. However, we observe a surprising failure mode: during instruction tuning, the context reliance under knowledge conflicts initially increases as expected, but then gradually decreases as instruction finetuning progresses. This happens while the performance on standard benchmarks keeps on increasing far after this drop. We call this phenomenon context-parametric inversion and observe it across multiple general purpose instruction tuning datasets such as TULU, Alpaca and Ultrachat, across different model families like Llama, Mistral, and Pythia. We perform various controlled studies and theoretical analysis to show that context-parametric inversion occurs due to examples in the instruction finetuning data where the input context provides information that aligns with model’s parametric knowledge. Our analysis suggests some natural mitigation strategies with limited but insightful gains, and serves as a useful starting point in addressing this deficiency in instruction finetuning.
@@ 1 Introduction
Large language models (LLMs) are widely used for a variety of tasks, many of which require effectively balancing the parametric knowledge embedded in their weights with the information provided through the input context. A persistent challenge, however, is their tendency to over-rely on this parametric knowledge, even when it contradicts the contextual information. This over-reliance not only hinders the ability to update model facts or mitigate hallucinations with augmented contexts, but also exacerbates issues like biases and failures to reliably follow user instructions—particularly when they involve the generation of atypical or unconventional text  (McKenzie et al., 2024; Qiu et al., 2023; Adlakha et al., 2024).
This tension between the context and the model’s parametric knowledge has been commonly studied under the moniker of knowledge conflicts, with existing works exploring various strategies like contrastive decoding or augmented finetuning to encourage models to adhere to the context  (Shi et al., 2023; Yuan et al., 2024; Longpre et al., 2022; Chen et al., 2022). However, this challenge still persists (see §B) and is unlikely to go away just with scale (McKenzie et al., 2024). We have limited understanding of the underlying dynamics that drive models to rely heavily on parametric knowledge despite contradictory contextual information.
In this work, we study the effect of instruction finetuning (IFT)—a staple part of the LLM pipeline—on the ability to override pretrained knowledge through the context. IFT seeks to enhance the model’s ability to assist with user queries or instructions. Oftentimes, these user queries and instructions are context-based where the context provides critical information needed to complete the task. For instance, an instruction “What is the total price of my trip to Hawaii?” operates on a context “Context: [Itinerary List]”, and an instruction “Rank these famous soccer players based on these scores” could contain a context like: “[Scores Table].” In these circumstances, instruction tuned models are responsible for appropriately leveraging the input context to answer the user query, instead of purely relying on parametric knowledge. In principle, instruction tuning should improve the ability of models to use context over parametric knowledge. However, we make an intriguing observation where in the presence of a conflict between parametric and context information, the model’s reliance on context initially increases as expected but surprisingly starts decreasing as IFT progresses.
Figure 1: (a) Context-Parametric Inversion In the presence of conflicts between parametric and context information, a model’s reliance on context first increases and then decreases during the process of instruction finetuning. (b) Instruction tuning data includes both context-critical examples, where the context is essential, and non-context-critical examples, where responses can be generated using either the context or the model’s parametric knowledge. It is this latter group of non-context-critical examples that leads to a decline in context reliance (§ 4.3).
@@
We measure the context reliance by designing inputs contexts that suggest a fictional answer to a user query different from facts memorized from the pretraining corpus (§ 3.2). We evaluate context reliance across the IFT trajectory on multiple IFT datasets—TULU, Alpaca or UltraChat and across multiple model families—Llama, Pythia and Mistral. Across all these settings, we see that context reliance initially increases and then decreases, a phenomenon we call context-parametric inversion. In fact, this drop begins in early timesteps of IFT, while the performance on standard benchmarks (e.g., MMLU, GSM8k, SQuAD) keeps on increasing far after this drop. For example, as shown in Figure 1(a), the context reliance of Llama2-7B (as measured on knowledge conflict datasets (§ 3.2)) increases from 30% to 60% initially with IFT. However, it start dropping as the finetuning progresses further, dipping to around $35\%$.
Why do we observe context-parametric inversion with instruction tuning? The initial increase is apriori expected, as instruction tuning datasets typically contains datapoints that require utilizing context to complete the instruction, as discussed above. We perform controlled experiments to understand the subsequent decrease which is unexpected and detrimental. First, we observe that reliance on parametric knowledge over context extends to facts beyond those potentially “memorized” during the instruction finetuning process. Hence, this drop in context reliance is not simply due to the model acquiring new facts. Common instruction tuning datasets typically contain some datapoints that are purely about recall of pretrained knowledge, and do not involve context-dependent instructions. Could the drop be attributed to the presence of such points? We curate the datasets to only include context-dependent points but still see a drop in context reliance after an initial increase.
We analyze this phenomenon theoretically in a simple setting and uncover the optimization dynamic that explains context-parametric inversion. We can partition a generic dataset containing context-dependent datapoints into two categories: (i) context-critical datapoints where context provides key information needed to answer a user query that the model does not know beforehand (Fig. 1(b)), and (ii) non-context-critical datapoints where the context is approximately redundant with model’s parametric knowledge. We formally define these categories in § 4.3.
@@
In the early stages of training, context-critical points tend to have higher loss and therefore dominate the gradient signal, driving the model to focus on the context. However, as training progresses, the loss on context-critical points decreases, and the non-context-critical points dominate the gradient. At later stages, the model leverages its parametric knowledge to further reduce the loss on these non-context-critical points, thus reducing the context reliance. We formalize this dynamic in a one-layer transformer under natural simplifying assumptions when finetuning on a data mixture consisting of context-critical and non-context-critical points.
Finally, our analysis naturally leads us to some mitigation strategies by data curation, data augmentation, and regularization. These strategies are able to partially alleviate the drop in deep networks on real-world datasets, showing that our theoretical insights do translate to practical settings. However, as we discuss in § 6, these mitigation strategies each have fundamental limitations and tradeoffs.
Overall, in this work, we uncover a broad failure in IFT, where the model begins to rely more on the parametric knowledge than the input context when they conflict, despite an initial jump in context reliance. To the best of our knowledge, we are the first to point out this deficiency with instruction tuned models. We also provide empirical and theoretical understanding of this observation along with some initial mitigation strategies that we hope would serve as a useful starting point to address the fundamental challenge of improving context reliance in language models.
@@ 2 Related Works
Language models are often exposed to user input instructions and accompanying context, which at times gives information or requests a behavior at odds with model’s prior from pretraining. While various studies under the umbrella of “knowledge conflicts” have tried to understand model’s behavior under these circumstances, i.e. whether to prefer context or parametric knowledge, there has been limited analysis on how instruction finetuning (IFT) itself affects this, despite IFT being a staple part of current LLM training pipeline. Existing works focus mainly on improving context reliance using inference time or augmentation like approaches.
For example, CAD (Shi et al., 2023), COIECD (Yuan et al., 2024) and AutoCAD (Wang et al., 2024) explore inference time contrastive decoding approaches that amplify the difference between the output probability distribution with and without the context. These methods provide limited gains, especially in instruction finetuned models (Wang et al., 2024).  Zhou et al. (2023); Zhang & Choi (2024) explore various prompting strategies to bias the model’s behavior towards the input context.  Jin et al. (2024b) tries to build a mechanistic interpretation. On the other hand,  Longpre et al. (2022); Fang et al. (2024); Neeman et al. (2022); Li et al. (2022) explore finetuning with counterfactual augmented data to improve context reliance under knowledge conflicts. However, in § 6, we show that counterfactual data augmentation cannot fix all types of context-parametric conflicts (e.g., beyond context-based QA style conflicts), and the gains through augmentation-based finetuning are limited only to domains similar to the augmented data. Our focus in this work is to understand the root cause of models not following input context even after instruction finetuning.
@@
Understanding the effect of instruction finetuning on knowledge conflicts is of high relevance for retrieval augmented generation (RAG), an important practical usecase of LLMs. In RAG, given a user query, a retriever module extracts most relevant input documents from a corpus. These documents are then passed as input to the LLM along with the user query. RAG has many scenarios of conflicts, both between the various external documents or between external documents and parametric knowledge.  Guu et al. (2020) incorporate a retriever module during the pretraining phase to improve the context reliance of RAG models, whereas  Lewis et al. (2021) incorporate a retriever during finetuning. In the case of conflicts between external documents,  Jin et al. (2024a); Kortukov et al. (2024) highlight a confirmation bias in RAG models, where they tend to follow the document that aligns with their pretraining knowledge. Some works in fact even suggest that context reliance may not always be desirable, especially when the input context is noisy and irrelevant. For example,  Zhang et al. (2024) propose a training procedure to instead increase the model’s tendency to answer using parametric knowledge when the input context might be noisy.
Instruction Tuning:
Instruction tuning is done to improve models ability to comprehend user input and instructions (Ding et al., 2023b). Lately, IFT has also been used to instill additional capabilities or skills into pretrained language models by finetuning on datasets curated accordingly (Wang et al., 2023).  Biderman et al. (2024); Wang et al. (2022); Kotha et al. (2024); Luo et al. (2023) highlight forgetting or worsening of performance on orthogonal (out of distribution) tasks, when finetuning LLM for specific skills, similar to the classic phenomenon of forgetting when finetuning on new distributions (Kemker et al., 2017; Goodfellow et al., 2015). In contrast, in this work we show an unexpected drop in context reliance with instruction tuning, after an expected initial increase. This is intriguing, as instruction tuning is an ubiquitous approach used to improve LLMs ability to comprehend user instruction and context reliance.
@@
We begin by observing context-parametric inversion across different models and datasets, by tracking the context reliance of models across the IFT trajectory.
Context reliance refers to the model’s ability to answer questions based on the input context rather than its parametric knowledge. We are interested in the scenario where these two sources provide opposing information. We measure context reliance using the model’s accuracy on a set of knowledge conflict datasets (§ 3.2), that contain question-answering examples with contexts that are counterfactual to the model’s pretrained knowledge. We measure accuracy by entailment. Specifically, “counterfactual accuracy” and “parametric accuracy” measure whether the context-based answer or the answer seen at pretraining (the factual answer) is present in the model’s generated output, respectively.
@@ 3.1 Experiment Setup
We experiment using three open source large language models—Llama2-7B, Pythia6.9B, and Mistral7B. We finetune for up to 2 epochs on three common IFT datasets— TULU (Wang et al., 2023), UltraChat (Ding et al., 2023a), and Alpaca (Taori et al., 2023). We track the progress of IFT based on the performance on four standard benchmarks: GSM8k (Cobbe et al., 2021) (math), MMLU (Hendrycks et al., 2021) (general fact recall), SQuAD (Rajpurkar et al., 2016) (reading comprehension), and ARC-Challenge (Clark et al., 2018) (reasoning). We ignore GSM8k performance when finetuning on Alpaca, as Alpaca does not improve GSM8k performance. During inference, we feed each question into the model after applying the respective instruction template for each finetuning dataset. We refer the reader to Appendix A.2 for additional details.
@@ 3.2 Knowledge Conflict Datasets
We carefully design three knowledge conflict datasets to get an accurate measure of model’s context reliance. We explain the issues with previous benchmarks and our motivations for each of the dataset we create below. All datasets are available at https://github.com/locuslab/context-parametric-inversion and on huggingface at locuslab/context_parametric_conflict. We refer the reader to Appendix A.5 for some examples.
@@
    Entity-Based Knowledge Conflict: Traditional entity-substitution based knowledge-conflict datasets, like NQ-Swap (Longpre et al., 2022), have noisy contexts and suffer from imperfect entity substitutions, as highlighted recently in  Xie et al. (2024). This happens because the entity substitution models (Honnibal & Montani, 2017) are not able to recognize and replace all the occurrences of factual answers in the input. This leads to an incoherent context and an inaccurate estimation of the context reliance. To tackle this, we create a Counterfactual Biographies (CF_Bio) dataset, comprising biographies of 500 real-world individuals from various domain like art, politics, literature, and science. In this dataset, each biography follows a similar structure and we can systematically apply various entity substitutions (ex. substituting names, contribution, etc.) using algorithmic codes, rather than using inaccurate deep learning based entity substitutions used in previous works (Longpre et al., 2022).
@@
    Coherent Counterfactual Contexts: Recently  Xie et al. (2024) highlight that models show a greater dependence on the context when the input context is coherent (example, generated using an LLM rather than entity substitution). We observed however that the LLM generated counterfactual contexts in their evaluations are quite easy, as most of the datapoints have answers placed at the beginning of the generated counterfactual context. Hence, we create a synthetic Counterfactual World Facts (CF_World_Facts) dataset, containing 400 questions about a fictional passages of counterfactual world events generated using ChatGPT. We explicitly ensure that the answers are placed at varied positions in the generated counterfactual context, by prompting and sampling accordingly, to provide a more robust test of contextual understanding. We refer the reader to Appendix A.5 for further details and examples.
@@
[…]